# Preference-Based Gradient Estimation for ML-Based Approximate Combinatorial Optimization

### v1

@@ 5.3 Decoding at Test Time
In other words, in the limit, the approximation algorithm guided by $\bm{s}$ always finds optimal solution $\bm{y}$. The proofs can be found in Appendix B.
@@ 6 Experiments
We validate our approach on two well-known combinatorial optimization problems: the travelling salesman problem (TSP) and the minimum $k$-cut problem. For both problems, we synthetically generate problem instances and establish baselines as reference.
Unweighted graphs NOIgen+
Method Single run Best out of 3 runs Single run Best out of 3 runs
@@
   (I-MLE) 3.47%$\pm$0.14 (402ms) 0.51%$\pm$0.06 (1.01s) 7.12%$\pm$0.35 (444ms) 0.39%$\pm$0.06 (1.10s)
   (PBGE, ours) 0.39%$\pm$0.05 (398ms) 0.06%$\pm$0.04 (1.01s) 0.57%$\pm$0.06 (439ms) 0.08%$\pm$0.05 (1.08s)
Table 1: Minimum $k$-cut optimality gaps on graphs with 100 nodes, using Karger-Stein as decoder. Mean $\pm$ standard deviation were calculated over five evaluation runs on the same model parameters. In the supervised and self-supervised rows, the parentheses indicate how the GNN was trained. In the columns labelled “Best out of 3 runs”, the Karger–Stein algorithm is run three times on the same GNN outputs, and the best result is used.
6.1 Problem Instance Generation
For minimum $k$-cut, we use the established graph generator NOIgen (Nagamochi et al., 1994). Since it relies on dramatically scaling down the weights of edges that are in the minimum $k$-cut in order to avoid trivial solutions, it makes it easy for a GNN to identify the correct edges. To make the graphs more challenging, we extend NOIgen to also use graph structure to avoid trivial solutions, which allows us to scale down edge weights less dramatically. We call this improved graph generator NOIgen+. We also generate unweighted graphs that only rely on graph structure to prevent trivial solutions. Graphs for the TSP are generated according to the established method described in Kool et al. (2019); Joshi et al. (2019, 2021). Please refer to appendix D for further details.
@@ 6.1 Problem Instance Generation
Figure 2: The number of graphs for which the minimum $k$-cut was found after a given number of Karger-Stein runs. For example, for Karger-Stein without a GNN on graphs with 100 nodes (2(a)), two Karger-Stein runs suffice to find the minimum $k$-cut for 83% of graphs in the validation set. The ideal algorithm, which always finds the minimum $k$-cut on the first attempt, would have simply a horizontal line at 100%.
6.2 Baselines
@@
and estimate $\nabla_{\bm{s}}\mathcal{L}(\mathcal{D},\bm{s})$ using REINFORCE (Williams, 1992) in the case of TSP and I-MLE (Niepert et al., 2021) in the case of the minimum $k$-cut problem. Additionally, for TSP, we also compare against a simple baseline that runs random insertion on the input graph 20 times and treats the best solution found as ground truth for a binary cross-entropy loss. We call this baseline “Best-of-20”.
For additional context, we train models in a supervised fashion using an edge-level binary cross-entropy loss comparing the GNN’s output scores $\bm{s}$ with a pre-calculated ground truth solution $\bm{y}$. For minimum $k$-cut, we also train using I-MLE, comparing the approximation algorithm’s output $\hat{\bm{y}}$ with $\bm{y}$ using a Hamming loss. See Appendix C for details on these baselines.
6.3 Minimum $k$-Cut
We evaluate our method on the minimum $k$-cut problem, using the Karger–Stein algorithm as a base. Table 1 shows optimality gaps of the unmodified Karger–Stein algorithm, as well as several versions of our method. Each version augments the Karger–Stein algorithm with a GNN, and they differ by how the GNN was trained. Note that when augmenting the Karger–Stein algorithm with a GNN trained with PBGE, the optimality gap improves by an order of magnitude. On top of this, even though it didn’t use any ground truth solutions during training, the GNN trained with PBGE comes close to matching the GNN trained supervised with a binary cross entropy loss.
In practice, the most important metric is the number of runs it takes for Karger–Stein to find the optimal $k$-cut. If this number is low, we can run Karger–Stein a small number of times and be reasonably certain that the minimum $k$-cut was found. Figure 2 shows for how many graphs the minimum $k$-cut is found in a set number of runs, comparing the unmodified Karger–Stein algorithm with two versions that were augmented using a GNN. On both datasets, the augmented Karger–Stein algorithm needs much fewer runs to find the minimum $k$-cut, almost always finding it on the first attempt. Again, the GNN trained self-supervised with PBGE comes close to matching supervised performance.
6.4 Travelling Salesman Problem (TSP)
Method (decoder in parentheses) $n=20$ $n=50$ $n=100$
@@ 6.4 Travelling Salesman Problem (TSP)
This work
   REINFORCE (random ins., 20 runs) 8.02%$\pm$0.07 (769ms) 23.94%$\pm$0.11 (11.62s) 52.61%$\pm$0.42 (1.38m)
   REINFORCE (random ins., 100 runs) 4.75%$\pm$0.03 (3.64s) 11.23%$\pm$0.08 (57.85s) 31.09%$\pm$0.26 (7.53m)
   Best-of-20 (random ins., 20 runs) 0.39%$\pm$0.03 (787ms) 11.97%$\pm$0.69 (11.03s)
   Best-of-20 (random ins., 100 runs) 0.10%$\pm$0.01 (3.70s) 5.28%$\pm$0.79 (52.82s)
   PBGE (ours) (random ins., 20 runs) 0.18%$\pm$0.01 (763ms) 2.36%$\pm$0.02 (11.11s) 5.15%$\pm$0.04 (1.43m)
   PBGE (ours) (random ins., 100 runs) 0.05%$\pm$0.01 (3.73s) 1.14%$\pm$0.03 (53.98s) 3.67%$\pm$0.05 (7.44m)
Khalil et al. (2017)
   (greedy) 1.42%* 5.16%* 7.03%*
Kool et al. (2019)
   (greedy) 0.34%* 1.76%* (2s) 4.53%* (6s)
   (sampling 1280 times) 0.08%* (5m) 0.52%* (24m) 2.26%* (1h)
Supervised
This work
   BCE loss (random insertion, 20 runs) 0.15%$\pm$0.01 (787ms) 0.94%$\pm$0.03 (10.92s) 2.86%$\pm$0.04 (1.47m)
   BCE loss (random insertion, 100 runs) 0.04%$\pm$0.00 (3.75s) 0.58%$\pm$0.02 (54.79s) 1.74%$\pm$0.03 (7.35m)
Joshi et al. (2019)
   (greedy) 0.60%* (6s) 3.10%* (55s) 8.38%* (6m)
   (beam search, beam width 1280) 0.10%* (20s) 0.26%* (2m) 2.11%* (10m)
Sun & Yang (2023)
   (greedy) 0.10%* 0.24%*
   (sampling 16 times) 0.00%* 0.00%*
@@
Exact solvers
Concorde (Applegate et al., 2006) 0.00%* (1m) 0.00%* (2m) 0.00%* (3m)
Gurobi 0.00%* (7s) 0.00%* (2m) 0.00%* (17m)
Table 2: TSP optimality gaps, with mean $\pm$ standard deviation, calculated over five evaluation runs on the same model parameters. Results marked with * are values obtained from the indicated papers and therefore do not include standard deviations.
Table 2 shows the optimality gaps of our approach and its variants. All of our models were trained using random insertion as the CO approximation algorithm. For PBGE, we sampled 10 solutions from $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ and 10 solutions from $h(\hat{\bm{y}}\mid{\mathcal{G}})$.
The decoder used at test time is listed after the name of the respective method in parentheses. Greedy search starts from an arbitrary node, and follows the edge with the highest score to select the next node. This process of greedily following the best edge is repeated until each node has been visited once. To ensure that the resulting tour is valid, edges that lead to nodes that have already been visited are excluded. Beam search also starts with an arbitrary node, then explores the $b$ edges with the highest scores. This gives us $b$ partial solutions. In each iteration, each partial solution is expanded at its last node, and out of the resulting paths, the $b$ best partial solutions are kept. As before, edges that would lead to invalid tours are ignored. The parameter $b$ is called the beam width, and beam search with $b=1$ corresponds to greedy search. Sampling simply refers to sampling multiple solutions and using the best one.
@@ 7 Conclusion
Acknowledgements
We thank Aneesh Barthakur, Matteo Palmas and Roman Freiberg for fruitful discussions. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Arman Mielke. The contributions of Thilo Strauss were carried out while employed with ETAS Research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Algorithms for Specific Combinatorial Optimization Problems
@@ A.1 Minimum $k$-Cut Problem
A.1.1 Karger’s Algorithm
@@
A.1.2 Karger–Stein Algorithm
@@
     return $\min\!\big{\{}\textsc{Karger-Stein}({\mathcal{G}}_{1}),\;\textsc{Karger-Stein}({\mathcal{G}}_{2})\big{\}}$ // Return the lower-weight cut
@@
A.2 Random Insertion for the Travelling Salesman Problem
@@
The algorithm is summarized in Algorithm 3
@@ B.1 Probability of Finding the Optimal Minimum $k$-cut
Let ${\mathcal{G}}=(V,E)$ be an undirected graph, $n=|V|$, and $m=|E|$.
During the first $i-1$ iterations of the modified Karger’s algorithm, some edges have been merged away. Let $E_{i}$ be the set of edges that are left at iteration $i$ (this means $E_{1}=E$). The probability that a specific edge $e$ is selected for contraction at iteration $i$ is
@@
Theorem B.1.
Let $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ be Karger’s algorithm steered by scores $\bm{s}$, and let $\bm{y}\in\{0,1\}^{|E|}$ be a minimum $k$-cut on ${\mathcal{G}}=(V,E)$. Then,
$\lim_{\sigma(\bm{s})\to\bm{y}}h(\bm{y}\mid{\mathcal{G}},\bm{s})=1.$
In other words, in the limit, Karger’s algorithm steered by $\bm{s}$ always finds optimal solution $\bm{y}$.
@@
The result can trivially be extended to the Karger–Stein algorithm, a variant of Karger’s algorithm, and to weighted graphs ${\mathcal{G}}(V,E,w)$.
B.2 Probability of Finding the Optimal TSP-Tour
Let ${\mathcal{G}}=(V,E,w)$ be an undirected graph. For a given TSP tour $\hat{\bm{y}}\in\{0,1\}^{|E|}$, let $J(\hat{\bm{y}})$ be the length of the tour:
$J(\hat{\bm{y}})=\sum_{e\in\hat{\bm{y}}}w(e)$
Let $h(\hat{\bm{y}}\mid{\mathcal{G}})$ be an insertion algorithm, such as random insertion (Karg & Thompson, 1964). Let $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ be that algorithm steered by the GNN output scores $\bm{s}$.
Theorem B.2.
Let $\bm{y}\in\{0,1\}^{|E|}$ be a minimal length TSP tour on graph ${\mathcal{G}}$. Then,

### v2

@@ 5.3 Theoretical analysis
More rigorous formulations of these theorems and their proofs can be found in Appendix B.
@@ 6 Experiments
We validate our approach on two well-known combinatorial optimization problems: the travelling salesman problem (TSP) and the minimum $k$-cut problem. For both problems, we synthetically generate problem instances and establish baselines as reference. We use residual gated graph convnets (Bresson & Laurent, 2017), but adapt them to include edge features $\bm{e}_{ij}^{l}$ and a dense attention map $\bm{\eta}_{ij}^{l}$ following Joshi et al. (2019). Please see Appendix C for details.
Table 1: Minimum $k$-cut optimality gaps on graphs with 100 nodes, using Karger–Stein as decoder. Mean $\pm$ standard deviation were calculated over ten evaluation runs on the same model parameters. In the supervised and self-supervised rows, a GNN trained with the indicated method is used to guide the Karger–Stein algorithm. In the columns labeled “Best out of 3 runs”, the Karger–Stein algorithm is run three times on the same GNN outputs, and the best result is used.
Method Unweighted graphs NOIgen+
Single run Best out of 3 runs Single run Best out of 3 runs
@@
   I-MLE 3.39%$\pm$0.16 (402ms) 0.52%$\pm$0.05 (1.01s) 7.63%$\pm$0.41 (444ms) 0.41%$\pm$0.07 (1.10s)
   PBGE (ours) 0.38%$\pm$0.05 (398ms) 0.06%$\pm$0.05 (1.01s) 0.58%$\pm$0.06 (439ms) 0.09%$\pm$0.05 (1.08s)
6.1 Problem instance generation and baselines
For minimum $k$-cut, we use the established graph generator NOIgen (Nagamochi et al., 1994). Since it relies on dramatically scaling down the weights of edges that are in the minimum $k$-cut in order to avoid trivial solutions, it makes it easy for a GNN to identify the correct edges. To make the graphs more challenging, we extend NOIgen to also use graph structure to avoid trivial solutions, which allows us to scale down edge weights less dramatically. We call this improved graph generator NOIgen+. We also generate unweighted graphs that only rely on graph structure to prevent trivial solutions. Graphs for the TSP are generated according to the established method described in Kool et al. (2019); Joshi et al. (2019, 2021). Please refer to Appendix E for further details.
@@ 6.1 Problem instance generation and baselines
Figure 2: The number of graphs for which the minimum $k$-cut was found after a given number of Karger–Stein runs. For example, for Karger–Stein without a GNN on graphs with 100 nodes (2(a)), two Karger–Stein runs suffice to find the minimum $k$-cut for 83% of graphs.
@@
and estimate $\nabla_{\bm{s}}\mathcal{L}(\mathcal{D},\bm{s})$ using I-MLE (Niepert et al., 2021), and, in the case of TSP, REINFORCE (Williams, 1992). Here, I-MLE acts as a representative of the decision-focused learning paradigm (Mandi et al., 2024).
For minimum $k$-cut, we also train supervised baselines using an edge-level binary cross entropy (BCE) loss comparing the GNN’s output scores $\bm{s}$ with a pre-calculated ground truth solution $\bm{y}$. Additionally, we train models using I-MLE in a supervised fashion, comparing the approximation algorithm’s output $\hat{\bm{y}}$ with $\bm{y}$ using a Hamming loss. See Appendix D for details on these baselines.
6.2 Results
Minimum $k$-Cut.
We evaluate our method on the minimum $k$-cut problem, using the Karger–Stein algorithm as a base. Table 1 shows optimality gaps of the unmodified Karger–Stein algorithm, as well as several versions of our method. Each version augments the Karger–Stein algorithm with a GNN, and they differ by how the GNN was trained. Note that when augmenting the Karger–Stein algorithm with a GNN trained with PBGE, the optimality gap improves by an order of magnitude. On top of this, even though it didn’t use any ground truth solutions during training, the GNN trained with PBGE comes close to matching the GNN trained supervised with a BCE loss.
In practice, the most important metric is the number of runs for Karger–Stein to find the optimal $k$-cut. If this number is low, we can run Karger–Stein a small number of times and be reasonably certain that the minimum $k$-cut was found. Figure 2 shows for how many graphs the minimum $k$-cut is found in a set number of runs, comparing the unmodified Karger–Stein algorithm with two versions that were augmented using a GNN. On both datasets, the augmented Karger–Stein algorithm needs much fewer runs to find the minimum $k$-cut, almost always finding it on the first attempt. Again, the GNN trained self-supervised with PBGE comes close to matching supervised performance.
Table 2: TSP optimality gaps, with mean $\pm$ standard deviation, calculated over ten evaluation runs on the same model parameters. We indicate the decoder used in parentheses. Values with * are obtained from these papers and therefore do not include standard deviations. We provide an extended version of this table with more baselines in Appendix F.
Method (decoder in parentheses) $\bm{n=20}$ $\bm{n=50}$ $\bm{n=100}$
@@ 6.2 Results
Bello et al. (greedy) 1.42%* 4.46%* 6.90%*
Khalil et al. (greedy) 1.42%* 5.16%* 7.03%*
Deudon et al. (greedy) 0.66%* (2m) 3.98%* (5m) 8.41%* (8m)
Deudon et al. (greedy + 2OPT) 0.42%* (4m) 2.77%* (26m) 5.21%* (3h)
Deudon et al. (sampling 1280 times) 0.11%* (5m) 1.28%* (17m) 12.70%* (56m)
Deudon et al. (sampling 1280 times + 2OPT) 0.09%* (6m) 1.00%* (32m) 4.64%* (5h)
Kool et al. (greedy) 0.34%* 1.76%* (2s) 4.53%* (6s)
Kool et al. (sampling 1280 times) 0.08%* (5m) 0.52%* (24m) 2.26%* (1h)
REINFORCE (random ins., 20 runs) 7.98%$\pm$0.08 (769ms) 23.84%$\pm$0.10 (11.62s) 52.83%$\pm$0.47 (1.38m)
REINFORCE (random ins., 100 runs) 4.78%$\pm$0.03 (3.64s) 11.24%$\pm$0.08 (57.85s) 31.06%$\pm$0.25 (7.53m)
I-MLE (random ins., 20 runs) 10.54%$\pm$0.15 (761ms) 35.87%$\pm$0.34 (11.14s) 61.42%$\pm$0.23 (1.37m)
I-MLE (random ins., 100 runs) 6.55%$\pm$0.10 (3.68s) 19.40%$\pm$0.24 (56.63s) 44.42%$\pm$0.14 (7.51m)
PBGE (ours) (random ins., 20 runs) 0.18%$\pm$0.01 (763ms) 2.37%$\pm$0.02 (11.11s) 5.12%$\pm$0.06 (1.43m)
PBGE (ours) (random ins., 100 runs) 0.05%$\pm$0.01 (3.73s) 1.13%$\pm$0.03 (53.98s) 3.67%$\pm$0.05 (7.44m)
@@
Figure 3: The Pareto frontier of self-supervised methods on TSP-100. Non-Pareto-optimal baselines with very high optimality gaps or very long runtimes are omitted for clarity.
Travelling salesman problem (TSP).
Table 2 shows the optimality gaps of our approach and its variants on TSP. All of our models were trained using random insertion as the CO approximation algorithm. For PBGE, we sampled 10 solutions from $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ and 10 solutions from $h(\hat{\bm{y}}\mid{\mathcal{G}})$.
The decoder used at test time is listed after the name of the respective method in parentheses. Greedy search starts from an arbitrary node, and follows the edge with the highest score to select the next node. This process of greedily following the best edge is repeated until each node has been visited once. To ensure that the resulting tour is valid, edges that lead to nodes that have already been visited are excluded. Sampling simply refers to sampling multiple solutions and using the best one. “+ 2OPT” refers to improving the decoded solution using 2OPT local search (Croes, 1958).
Figure 3 shows the Pareto frontier of self-supervised methods on TSP-100. PBGE achieves Pareto-optimality when decoding with 100 random insertion runs.
@@ 7 Conclusion
Acknowledgments and Disclosure of Funding
The authors would like to thank Aneesh Barthakur, Matteo Palmas, Roman Freiberg, and Jiaqi Wang for fruitful discussions. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Arman Mielke. The contributions of Thilo Strauss were carried out while employed with ETAS Research.
Appendix A Combinatorial optimization problems and approximation algorithms
@@ A.1 Minimum $k$-cut problem
A.1.1 Problem Definition
We are given a connected, undirected graph ${\mathcal{G}}=(V,E,w)$ with edge weights $w:E\rightarrow\mathbb{R}_{>0}$, as well as a desired number of connected components $k\in\mathbb{N},2\leq k\leq|V|$. The goal is to find a set of edges $C\subseteq E$ with minimal total weight whose removal leaves $k$ connected components. This set is called a minimum $k$-cut. Formally, we are optimizing
$\min_{C\subseteq E}\sum_{e\in C}w(e)\quad\text{ such that graph }(V,E\backslash C)\text{ has }k\text{ connected components}.$
A.1.2 Karger’s algorithm
@@
A.1.3 Karger–Stein algorithm
@@
     return $\min\!\big{\{}\textsc{Karger--Stein}({\mathcal{G}}_{1}),\;\textsc{Karger--Stein}({\mathcal{G}}_{2})\big{\}}$triangleright Return the lower-weight cut
@@
A.2 Travelling Salesman Problem (TSP)
A.2.1 Problem Definition
We are given a complete directed or undirected graph ${\mathcal{G}}=(V,E,w)$ with edge weights $w:E\longrightarrow\mathbb{R}$. Our goal is to find a minimum weight Hamiltonian cycle in ${\mathcal{G}}$. A Hamiltonian cycle, also called a tour, is a sequence of nodes where each node in the graph appears exactly once, each node in the sequence is adjacent to the previous node, and the first node is adjacent to the last node. The weight of the cycle is the sum of all edge weights of edges between nodes that appear next to each other in the cycle, including the edge from the last to the first node. Instead of representing a tour as a sequence of nodes, we will represent it as the set of edges between nodes that appear next to each other in the sequence. Formally, we are optimizing
$\min_{T\subseteq E}\sum_{e\in T}w(e)\quad\text{ such that }T\text{ forms a Hamiltonian cycle}.$
The nodes are commonly called cities, and the weight of an edge is commonly called the distance between the two cities. The TSP is NP complete.
Metric TSP.
The metric version of TSP additionally assumes that the distances between the cities form a metric. This means that
  • the graph is undirected (or $w(x,y)=w(y,x)$ for all $x,y\in V$),
  • distances between cities at different locations are positive, and
  • the edge weights satisfy the triangle inequality, i.e. $w(x,y)+w(y,z)\geq w(x,z)$ for all $x,y,z\in V$.
Euclidean TSP.
Euclidean TSP is a special case of metric TSP in which the cities are located at points in the unit square and the distances between the cities are the Euclidean distances between the respective points. Research on solving the TSP using neural networks often focuses on this version of the problem (Kool et al., 2019; Joshi et al., 2019, 2021).
A.2.2 Random insertion algorithm
@@
The algorithm is summarized in Algorithm 3.
@@ B.1 Probability of finding the optimal minimum $k$-cut
Let ${\mathcal{G}}=(V,E)$ be an undirected graph. During the first $i-1$ iterations of the modified Karger’s algorithm, some edges have been merged away. Let $E_{i}$ be the set of edges that are left at iteration $i$ (this means $E_{1}=E$). The probability that a specific edge $e$ is selected for contraction at iteration $i$ is
@@
Theorem 1.
Let ${\mathcal{G}}=(V,E)$ be an undirected graph, and let $\bm{y}\in\{0,1\}^{|E|}$ be a minimum $k$-cut on ${\mathcal{G}}$. Moreover, let $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ be Karger’s algorithm guided by parameters $\bm{s}\in\mathbb{R}^{|E|}$, and let $\sigma$ be the element-wise sigmoid function. Then, $\lim_{\sigma(\bm{s})\to\bm{y}}h(\bm{y}\mid{\mathcal{G}},\bm{s})=1.$
In other words, in the limit, Karger’s algorithm guided by $\bm{s}$ always finds optimal solution $\bm{y}$.
@@
The result can trivially be extended to the Karger–Stein algorithm, a variant of Karger’s algorithm, and to weighted graphs ${\mathcal{G}}=(V,E,w)$.
B.2 Probability of finding optimal solutions using algorithms with relative error bounds
Let ${\mathcal{G}}=(V,E,w)$ be a weighted graph with edge weights $w:E\rightarrow\mathbb{R}_{>0}$. Consider any CO problem that asks to find a set of edges of minimum total weight that fits the constraints (e.g. minimum $k$-cut or TSP). This means that the CO problem’s objective function takes the form $J(\hat{\bm{y}})=\sum_{e\in\hat{\bm{y}}}w(e)$. Remember that we allow ourselves the notational convenience of treating $\hat{\bm{y}}$ as both a set of edges and a vector in $\{0,1\}^{|E|}$.
Theorem 2.
Let $J$ be the objective of a minimization problem on graphs with $J(\hat{\bm{y}})=\sum_{e\in\hat{\bm{y}}}w(e)$ and $w(e)>0$. Let ${\mathcal{G}}=(V,E,w)$ be an undirected, weighted graph, and let $\bm{y}\in\{0,1\}^{|E|}$ be an optimal solution on ${\mathcal{G}}$. Moreover, let $h(\hat{\bm{y}}\mid{\mathcal{G}})$ be a probabilistic approximation algorithm with a bound on the relative error, that is, there exists a function $\alpha$ such that for all $\hat{\bm{y}}\sim h(\hat{\bm{y}}\mid{\mathcal{G}})$, $J(\hat{\bm{y}})\leq\alpha({\mathcal{G}})J(\bm{y})$. Finally, let $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ be the parameterized version of that approximation algorithm, using parameters $\bm{s}\in\mathbb{R}^{|E|}$, and let $\sigma$ be the element-wise sigmoid function. Then, $\lim_{\sigma(\bm{s})\to\bm{y}}h(\bm{y}\mid{\mathcal{G}},\bm{s})=1.$
Proof.
Let ${\mathcal{G}}^{\prime}$ be the graph that the approximation algorithm receives as input after modification through $\bm{s}$, i.e. $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})=h(\hat{\bm{y}}\mid{\mathcal{G}}^{\prime})$. Let $J_{{\mathcal{G}}^{\prime}}(\bm{y})$ denote the total weight of a solution in the modified graph ${\mathcal{G}}^{\prime}$. As $\sigma(\bm{s})$ approaches $\bm{y}$, the weights in ${\mathcal{G}}^{\prime}$ of edges in $\bm{y}$ approach 0. This means that the total weight of the optimal solution $\bm{y}$ in the modified graph ${\mathcal{G}}^{\prime}$ approaches 0:
$\lim_{\sigma(\bm{s})\to\bm{y}}J_{{\mathcal{G}}^{\prime}}(\bm{y})=0.$
Any other solution $\bm{y}^{\prime}$ cannot be a strict subset of the edges included in $\bm{y}$, since that would mean that it has a lower total weight than $\bm{y}$. $\bm{y}^{\prime}$ therefore includes at least one edge that is not included in $\bm{y}$. Since this edge’s original edge weight is greater than 0, its scaled weight remains greater than 0 as $\sigma(\bm{s})$ approaches $\bm{y}$. The total weight in ${\mathcal{G}}^{\prime}$ of any solution other than $\bm{y}$ therefore approaches a value greater than 0:
$\lim_{\sigma(\bm{s})\to\bm{y}}J_{{\mathcal{G}}^{\prime}}(\bm{y}^{\prime})>0\quad\forall\bm{y}^{\prime}\neq\bm{y}.$
In the limit, the right hand side of the relative error bound $J_{{\mathcal{G}}^{\prime}}(\hat{\bm{y}})\leq\alpha(G)J_{{\mathcal{G}}^{\prime}}(\bm{y})$ approaches 0. This means that the maximum total weight of solutions that the approximation algorithm can find approaches 0, and, as soon as the lengths of the suboptimal solutions in ${\mathcal{G}}^{\prime}$ are far enough away from 0, $\bm{y}$ is the only solution left that the approximation algorithm can find. ∎
Note that this proof can trivially be adapted to CO problems that asks to find a set of nodes of minimum total weight that fits the constraints.
We know that the TSP tour produced by any insertion algorithm $h(\hat{\bm{y}}\mid{\mathcal{G}})$, such as random insertion (Karg & Thompson, 1964), is within $\lceil\log(|V|)\rceil+1$ of the optimal tour (Rosenkrantz et al., 1977). Formally, this means that $J(\hat{\bm{y}})\leq\big{(}\lceil\log(|V|)\rceil+1\big{)}\>J(\bm{y})$ for $\hat{\bm{y}}\sim h(\hat{\bm{y}}\mid{\mathcal{G}})$. From that, the following result immediately follows:
Corollary 3.
Let ${\mathcal{G}}=(V,E,w)$ be an undirected, weighted graph, and let $\bm{y}\in\{0,1\}^{|E|}$ be a minimal length TSP tour on ${\mathcal{G}}$. Let $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ be a probabilistic insertion algorithm, guided by parameters $\bm{s}\in\mathbb{R}^{|E|}$. Let $\sigma$ be the element-wise sigmoid function. Then,